# Intrinsic Heisenberg-type lower bounds on spacelike hypersurfaces in general relativity

### v1

@@ Abstract
We prove a coordinate- and foliation-independent Heisenberg-type lower bound for quantum states strictly localized in geodesic balls of radius $r$ on spacelike hypersurfaces of arbitrary spacetimes (with matter and a cosmological constant). The estimate depends only on the induced Riemannian geometry of the slice; it is independent of the lapse, shift, and extrinsic curvature, and controls the canonical momentum variance/uncertainty $\sigma_{p}$ by the first Dirichlet eigenvalue of the Laplace–Beltrami operator (Theorem). On weakly mean-convex balls we obtain the universal product inequality $\sigma_{p}\,r\geq\hbar/2$, whose constant is optimal and never attained (Corollary). This result abstracts and extends the framework recently introduced for black-hole slices in [1].
@@ 1 Introduction
On a Riemannian manifold $(\Sigma,h)$, strict localization of a quantum state to a bounded domain $\Omega\subset\Sigma$ with Dirichlet boundary conditions entails a non-vanishing spectral cost for the momentum observable: once $\psi|_{\partial\Omega}=0$, the $L^{2}$-norm of its gradient cannot be arbitrarily small. In spectral terms, the intrinsic uncertainty is governed by the first Dirichlet eigenvalue of the Laplace–Beltrami operator on $\Omega$ [2, Ch. 3]. This paper applies this principle to canonical data on Cauchy slices of curved spacetimes, making it directly usable from a physics perspective. On any spacelike hypersurface $(\Sigma,h)$ in a curved spacetime, we prove the following lower bound for states strictly localized to geodesic balls $B_{\Sigma}(p,r)$ [^1] Here $\lambda_{1}$ denotes the first Dirichlet eigenvalue of the Laplace–Beltrami operator $\Delta_{h}$ on $B_{\Sigma}(p,r)$.
@@
Beyond the main bound, we obtain a universal Heisenberg-type inequality
$\sigma_{p}(\psi)\,r\geq\frac{\hbar}{2}$
on weakly mean-convex geodesic balls of radius $r$, see Corollary 3.6. The argument relies solely on the boundary-distance Hardy inequality, available under the distributional superharmonicity of the distance to $\partial B_{\Sigma}(p,r)$ [3, 4, 5]. This baseline is never attained and its constant is optimal, hence it can be used as a robust, geometry-free floor whenever only coarse boundary information is accessible.
The results here abstract and extend the framework introduced for black-hole slices in [1]: no symmetry, stationarity, or vacuum assumption is required, and the statement is purely Riemannian on the chosen slice. When curvature information is available, classical comparison techniques immediately transfer to quantitative control of $\lambda_{1}$ and thus of $\sigma_{p}$ : see, e.g., Cheng-type eigenvalue comparison theorems [6, 7], small-ball asymptotics [8], and explicit radial analyses on spherically symmetric manifolds [9]. In constant positive curvature, our universal product floor coexists with stronger model-specific bounds derived in [10].
Organization. Section 2 introduces the minimal geometric setting and the precise Dirichlet localization; in particular, Assumption 2.1 records the standing conditions for $B_{\Sigma}(p,r)$. Section 3 establishes the intrinsic Heisenberg bound (Theorem 3.1). The subsection containing Corollary 3.6 derives the universal Hardy-Heisenberg baseline (19). Section 4 closes with a discussion of scope, extensions (including spinors/forms), and computational aspects.
@@ 3 An intrinsic Heisenberg-type lower bound
By the Dirichlet Rayleigh–Ritz characterization (6) (see [2, Ch. 3]), for any $0\neq u\in H^{1}_{0}(B_{\Sigma}(p,r))$,
@@ 3.1 Universal Hardy baseline on weakly mean-convex balls
Let $(\Sigma,h)$ be a Riemannian manifold and $\Omega\subset\Sigma$ a bounded $C^{2}$ domain. We say that $\Omega$ is weakly mean-convex if the mean curvature $H$ of $\partial\Omega$ with respect to the outward unit normal is nonnegative in the distributional sense. Equivalently, the boundary-distance function $d(x)=\operatorname{dist}_{h}(x,\partial\Omega)$ is distributionally superharmonic on $\Omega$ (i.e., $\Delta_{h}d\geq 0$) – see [3, Thm. 1.2] and [4].
@@
is valid on the larger interval $0<r<\pi/\sqrt{\kappa}$ and decreases to 0 as $r\uparrow\pi/\sqrt{\kappa}$ (antipodal radius), where mean-convexity fails. On the overlap $0<r\leq\pi/(2\sqrt{\kappa})$ the bound (22) is strictly stronger than (19); for instance, at $r=\pi/(2\sqrt{\kappa})$ it gives $\sigma_{p}\,r\geq\pi\hbar\sqrt{3}/2\gg\hbar/2$. Thus there is no contradiction: (19) is a foliation-independent, boundary-driven floor available under mean-convexity, while (22) is a curvature-sensitive refinement that extends beyond the convexity radius.
@@ 4 Discussion, scope, and extensions
Beyond the Hardy floor, curvature hypotheses convert geometric control into spectral control of $\lambda_{1}$ and hence into quantitative control of $\sigma_{p}$. Classical eigenvalue comparison theorems under sectional or Ricci curvature bounds (Cheng-type inequalities) yield effective upper and lower bounds for the first Dirichlet eigenvalue of geodesic balls; see [6, 7]. In the small-radius regime, asymptotic expansions quantify the departure of $\lambda_{1}$ from its Euclidean value in terms of curvature at the center point [8]. On spherically symmetric manifolds, the radial structure can be exploited to compute or bound spectra explicitly [9]. Each of these inputs transfers directly to $\sigma_{p}$ via Theorem 3.1.
@@
Taken together, the universal Hardy baseline and curvature-sensitive refinements yield a two-tier structure: a geometry-free minimum available under weak mean-convexity and, when curvature information is present, sharper controls tailored to the ambient geometry. This template extends naturally to other strictly local observables, to anisotropic comparison results adapted to symmetry classes or foliations, and to measurement-theoretic formulations of quantum mechanics on curved backgrounds.

### v2

@@ Abstract
We prove a coordinate- and foliation-independent Heisenberg-type lower bound for quantum states strictly localized in geodesic balls of radius $r$ on spacelike hypersurfaces of arbitrary spacetimes (with matter and a cosmological constant). The estimate depends only on the induced Riemannian geometry of the slice; it is independent of the lapse, shift, and extrinsic curvature, and controls the canonical momentum variance/uncertainty $\sigma_{p}$ by the first Dirichlet eigenvalue of the Laplace–Beltrami operator (Theorem). On weakly mean-convex balls we obtain the universal product inequality $\sigma_{p}r\geq\hbar/2$. Under the same assumption, a vector-field Barta-type argument improves this universal floor to the scale-invariant bound $\sigma_{p}r\geq\pi\hbar/2$, which provides a universal, foliation-independent floor. Any further sharpening of the constant requires eigenvalue‑comparison results or other curvature‑sensitive methods.
@@ 1 Introduction
On a Riemannian manifold $(\Sigma,h)$, strict localization of a quantum state to a bounded domain $\Omega\subset\Sigma$ with Dirichlet boundary conditions entails a non-vanishing spectral cost for the momentum observable: once $\psi|_{\partial\Omega}=0$, the $L^{2}$-norm of its gradient cannot be arbitrarily small. In spectral terms, the intrinsic uncertainty is governed by the first Dirichlet eigenvalue of the Laplace–Beltrami operator on $\Omega$ [1, Ch. 3]. This paper applies this principle to canonical data on Cauchy slices of curved spacetimes, making it directly usable from a physics perspective. On any spacelike hypersurface $(\Sigma,h)$ in a curved spacetime, we prove the following lower bound for states strictly localized to geodesic balls $B_{\Sigma}(p,r)$ [^1] Here $\lambda_{1}$ denotes the first Dirichlet eigenvalue of the Laplace–Beltrami operator $\Delta_{h}$ on $B_{\Sigma}(p,r)$.
@@
Beyond the main bound, we obtain a universal Heisenberg-type product inequality $\sigma_{p}\,r\geq\hbar/2$}, on weakly mean-convex geodesic balls of radius $r$; see Corollary 3.6. The argument uses only the boundary-distance Hardy inequality, valid under the distributional superharmonicity of the distance to $\partial B_{\Sigma}(p,r)$ [2, 3, 4]. The constant $1/2$ is optimal and the bound is never attained, so it serves as a robust, geometry-free floor whenever only coarse boundary information is available.
Under the same weak mean-convexity hypothesis, Corollary 3.7 sharpens this baseline by a factor of $\pi$ via a vector-field version of Barta’s method, yielding the scale-invariant estimate (23) [5]. This improvement depends only on the sign of the radial Laplacian and remains coordinate- and foliation-independent within the present framework; in particular, no symmetry, stationarity, or vacuum assumption is required on the chosen slice.
When curvature information is available, classical comparison techniques immediately transfer to quantitative control of $\lambda_{1}$ and thus of $\sigma_{p}$ : see, e.g., Cheng-type eigenvalue comparison theorems [6, 7], small-ball asymptotics [8], and explicit radial analyses on spherically symmetric manifolds [9]. In constant positive curvature, our universal product floor coexists with stronger model-specific bounds derived in [10].
Organization. Section 2 introduces the minimal geometric setting and the precise Dirichlet localization; in particular, Assumption 2.1 records the standing conditions for $B_{\Sigma}(p,r)$. Section 3 establishes the intrinsic Heisenberg bound (Theorem 3.1). The subsections 3.1 and 3.2 derive the universal Hardy-Heisenberg baseline and the Barta-type improvement. Section 4 closes with a discussion of scope, extensions (including spinors/forms), and computational aspects.
@@ 3 An intrinsic Heisenberg-type lower bound
By the Dirichlet Rayleigh–Ritz characterization (6) (see [1, Ch. 3]), for any $0\neq u\in H^{1}_{0}(B_{\Sigma}(p,r))$,
@@ 3.1 Universal Hardy baseline on weakly mean-convex balls
Let $(\Sigma,h)$ be a Riemannian manifold and $\Omega\subset\Sigma$ a bounded $C^{2}$ domain. We say that $\Omega$ is weakly mean-convex if the mean curvature $H$ of $\partial\Omega$ with respect to the outward unit normal is nonnegative in the distributional sense. Equivalently, the boundary-distance function $d(x)=\operatorname{dist}_{h}(x,\partial\Omega)$ is distributionally superharmonic on $\Omega$ (i.e., $\Delta_{h}d\geq 0$) – see [2, Thm. 1.2] and [3].
@@
is valid on the larger interval $0<r<\pi/\sqrt{\kappa}$ and decreases to 0 as $r\uparrow\pi/\sqrt{\kappa}$ (antipodal radius), where mean-convexity fails. On the overlap $0<r\leq\pi/(2\sqrt{\kappa})$ the bound (22) is strictly stronger than (19); for instance, at $r=\pi/(2\sqrt{\kappa})$ it yields $\sigma_{p}\,r\geq\pi\hbar\sqrt{3}/2\gg\hbar/2$. Thus there is no contradiction: (19) is a foliation-independent, boundary-driven floor available under mean-convexity, while (22) is a curvature-sensitive refinement that extends beyond the convexity radius.
3.2 Barta-type improvement under weak mean-convexity
The universal Hardy baseline already enforces the scale-invariant floor (19) under the weak mean-convexity hypothesis in Assumption 2.1. In this subsection we sharpen that floor using a vector-field version of Barta’s method [5]. The idea is entirely intrinsic to $(\Sigma,h)$: one tests the Dirichlet form with a radially aligned drift field adapted to the distance function, so that only the sign of the radial Laplacian enters. Combined with the intrinsic lower bound of Theorem 3.1, this yields a factor-$\pi$ improvement of the product inequality, recorded as (23). The proof is short and foliation-independent, requires no symmetry or vacuum assumptions, and rests on the Barta identity in the formulation of [5]. We state the result next as Corollary 3.7.
Corollary 3.7 (Barta-type improvement under weak mean-convexity).
Assume that $B=B_{\Sigma}(p,r)$} is weakly mean-convex, equivalently $\Delta\rho\geq 0$ on $B\!\setminus\!\{p\}$, where $\rho(x):=\mathrm{dist}_{h}(x,p)$. Then
$\sigma_{p}(\psi)\,r\ \geq\ \frac{\pi\hbar}{2}\,$ (23)
for all $\psi\in H^{1}_{0}(B),\|\psi\|_{2}=1$.
Proof.
We use the vector-field version of Barta’s method from [5]. For any $X\in C^{1}(B,T\Sigma)$ and any $u\in H^{1}_{0}(B)$ one has the identity
$\int_{B}|\nabla u|^{2}=\int_{B}\bigl|\nabla u-Xu\bigr|^{2}-\int_{B}(\mathrm{div}X+|X|^{2})\,u^{2},$ (24)
hence
$\int_{B}|\nabla u|^{2}\ \geq\ \int_{B}\bigl(-\mathrm{div}X-|X|^{2}\bigr)\,u^{2}\quad\Longrightarrow\quad\lambda_{1}(B)\ \geq\ \inf_{B}\bigl(-\mathrm{div}X-|X|^{2}\bigr).$ (25)
Choose $\alpha:=\pi/(2r)$ and the radial field
$X\ :=\ -\,\alpha\,\tan\bigl(\alpha\,\rho\bigr)\,\nabla\rho.$
Using $\mathrm{div}(\phi(\rho)\nabla\rho)=\phi^{\prime}(\rho)+\phi(\rho)\,\Delta\rho$ (valid on $B\setminus\{p\}$; a standard cut-off around $p$ justifies integration), and $|\nabla\rho|=1$, we compute
$\mathrm{div}X-|X|^{2}=\alpha^{2}+\alpha\,(\Delta\rho)\,\tan(\alpha\rho).$
Since $0\leq\alpha\rho<\tfrac{\pi}{2}$ on $B$ and $\tan(\alpha\rho)\geq 0$, the weak mean-convexity hypothesis $\Delta\rho\geq 0$ yields
$\mathrm{div}X-|X|^{2}\ \geq\ \alpha^{2}\ =\ \frac{\pi^{2}}{4r^{2}}.$
Taking the infimum over $B$ gives $\lambda_{1}(B)\geq\pi^{2}/(4r^{2})$. The second claim then follows from Theorem 3.1. ∎
Remark 3.8 (Context for the bounds (19) and (23)).
Under Assumption 2.1, the identity $d=r-\rho$ reduces the boundary-distance Hardy hypothesis to the radial condition used in the Barta argument. Thus the improvement from the universal Hardy baseline (19) to the Barta-type estimate (23) comes from the vector-field method and choice of $X$, not from a stronger geometric assumption. This improvement is scale-invariant and uses only the sign of $\Delta\rho$, hence it is robust, geometry- and foliation-independent within the present framework. In minimax sense and three dimensions, (23) is best possible under exactly the hypotheses of Corollary 3.7; any further strengthening requires additional information (e.g. curvature/comparison hypotheses), cf. [5, 6] or model-specific refinement [10].
@@ 4 Discussion, scope, and extensions
Beyond the Hardy floor a factor-$\pi$ improvement is obtained via the vector-field version of Barta’s method [5]. Together with Theorem 3.1, this yields a strictly stronger, model-agnostic control on weakly mean-convex balls, and it can only be further sharpened wherever eigenvalue comparison and curvature-sensitive techniques are available [6, 7].
Classical eigenvalue comparison theorems under sectional or Ricci curvature bounds (Cheng-type inequalities) yield effective upper and lower bounds for the first Dirichlet eigenvalue of geodesic balls; see [6, 7]. In the small-radius regime, asymptotic expansions quantify the departure of $\lambda_{1}$ from its Euclidean value in terms of curvature at the center point [8]. On spherically symmetric manifolds, the radial structure can be exploited to compute or bound spectra explicitly [9]. Each of these inputs transfers directly to $\sigma_{p}$ via Theorem 3.1.
@@
Taken together, the universal Hardy/Barta baseline and curvature-sensitive refinements yield a two-tier structure: a geometry-free minimum available under weak mean-convexity and, when curvature information is present, sharper controls tailored to the ambient geometry. This template extends naturally to other strictly local observables, to anisotropic comparison results adapted to symmetry classes or foliations, and to measurement-theoretic formulations of quantum mechanics on curved backgrounds.